OpenStudy (anonymous):

find the coordinates of the midpoint of hx?? H(5 1/2,-4 3/4) X(2 1/4, -1 1/4)

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ H&({\color{red}{ 5\frac{1}{2}}}\quad ,&{\color{blue}{ -4\frac{3}{4}}})\quad X&({\color{red}{ 2\frac{1}{4}}}\quad ,&{\color{blue}{ -1\frac{1}{4}}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)